Assume that X has a normal distribution, and find the indicated probability.
The mean is μ = 60.0 and the standard deviation is σ = 4.0.
Find the probability that X is less than 53.0.

Answer by stanbon(75887)

The mean is μ = 60.0 and the standard deviation is σ = 4.0.
Find the probability that X is less than 53.0.
--------
z(53) = (53-60)/4 = -7/4
-----
P(x < 53) = P(z < -7/4) = normalcdf(-100,-7/4) = 0.0401
=========================
Cheers,
Stan H.
=========================